A wire of length L and cross sectional area A is made of a material of Young’s modulus Y. If the wire is stretched by an amount x, the work done is………..

Grade:12th pass

1 Answers

Aditi Chauhan
9 years ago

W = 1/2 *Y*(strain)2* Yd

= 1/2 *Y* x2/L2*AL = YAx2/2L